A ball is thrown up with an initial speed of 20 ft/s. How high will it rise? How long will it take the body to reach the maximum height? How long will it be in air? Solve for the average speed
1
Expert's answer
2020-11-16T07:48:49-0500
"t=\frac{v}{g}=\frac{20\cdot0.3048}{9.8}=0.622\ s"

"v^2=2gh\\(20\cdot0.3048)^2=2(9.8)h\\h=1.9\ m"